OpenStudy (anonymous):

completing the square of 4x (3x+2)= x+12

OpenStudy (nurali):

4x (3x+2)= x+12 12x^2+8x=x+12 12x^2+8x-x-12=0 12x^2+7x-12=0 12x+16-9x-12=0 4x(3x+4)-3(3x+4)=0 (4x-3)(3x+4) = 0 x = 3/4 or x = -4/3

OpenStudy (nurali):

Welcome to OpenStudy @Uly

OpenStudy (whpalmer4):

I wouldn't call that completing the square, although it does get the right answer. \[4x(3x+2) = x+12\]\[12x^2+8x = x + 12\]\[12x^2+7x = 12\]Divide through by 12 to get leading coefficient of 1 for completing the square \[x^2+\frac{7}{12}x = 1\]Take half of the coefficient of \(x\), square, and add to both sides to complete the square \[x^2 + \frac{7}{12}x + (\frac{7}{24})^2 = 1 + (\frac{7}{24})^2\]Rewrite LHS as perfect square\[\frac{1}{24^2}(24x+7)^2 = \frac{625}{576}\]Cancel common factor \[(24x+7)^2 = 625\]Take square root of both sides\[24x+7 = \pm 25\]Solve for \(x\)\[24x = -7\pm25\]\[x = \frac{3}{4}, ~x = -\frac{4}{3}\]

OpenStudy (anonymous):

yeahh..totally..the first one is not a completing the square.